OpenStudy (anonymous):

The hypotenuse of a right triangle is 20 centimeters, and the shorter leg is 12 centimeters. Find the length of the other leg.

OpenStudy (anonymous):

12^2 + x^2 = 20^2

OpenStudy (anonymous):

144 + x^2 = 400

OpenStudy (anonymous):

x^2 = 400 - 144

OpenStudy (anonymous):

x^2 = 256

OpenStudy (anonymous):

Answer choices 2cm 4cm 16 cm forgot to type them onn their

OpenStudy (anonymous):

x = square root of 256

OpenStudy (anonymous):

x = 16